Pleas help me~
The sum of two numbers is 21. Find both numbers if twice the smaller number is 3 more than the larger number.
Thank you.

Answer by pie90010(44)

Let x be one # (the smaller one)
Let y be another # (the larger one)
x + y = 21
2x = y + 3
The first equation is fine, line up the second one
2x - y = 3
x + y = 21
the Y's cancel , now add them up
3x = 24
divide both sides by 3 and
x = 8
Now to get Y just plug it in to one of the original equations
8 + y = 21
y = 13



The two numbers are 8, 13